OpenStudy (joachim):

I really need some help with this question: "Two runners start simultaneously from the same point on a circular 200 m track and run in opposite directions. One runs at a constant speed of 6.20 m/s, and the other runs at a constant speed of 5.50 m/s. When they first meet, a) for how long a time will they have been running b) how far will each one have run along the track?" If you could just help with part a, i'll be able to get part b on my own. Thanks in advance :)

OpenStudy (anonymous):

The total distance they would have traveled together when they meet is equal to 200m. Therefore, one runner would have run (x)m while the other would have run (200 - x)m. You can then find the time taken. Time = distance / speed. Time taken would be the same for both. So, time taken = x/6.20 = (200 - x) / 5.50 Solve for x first then find the time taken.

OpenStudy (anonymous):

6,2x+5,5x=200 , which is 11,7x=200 , devide by 11.7 on both sides to get x, which is number of seconds